OpenStudy (anonymous):

(3.83)^2 + 7.6 x 5.7 (round your answer to the nearest hour) (show steps please and thank you!)

OpenStudy (anonymous):

3.83*3.83 = 14.669 7.6*5.7 =43.32 14.669+43.32 = 57.9889 round 57.9889 to 58 i could be wrong but im pretty sure that is correct

OpenStudy (anonymous):

to the nearest hour? what units are your numbers representing?

OpenStudy (anonymous):

Just round the answer to the nearest 10.